solve each system by elimination
4x-10y=6
8x-20y=24

Answer by ewatrrr(24785)

 
Hi
solve each system by elimination
4x-10y= 6 ||multiplying thru by -2 to eliminate the y variable
8x-20y= 24
-8x+20y= -12 ||result of multiplying thru by -2 to eliminate the y variable
8x-20y= 24
0 ≠ 12 there is NO solution for this System